Hi, Would you...Please help me.

Lupo Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on the following data: Total machine-hours 30,900 Total fixed manufacturing overhead cost $ 154,500 Variable manufacturing overhead per machine-hour $ 3.00 Recently, Job T687 was completed with the following characteristics: Number of units in the job 10 Total machine-hours 30 Direct materials $ 740 Direct labor cost $ 1,480 The amount of overhead applied to Job T687 is closest to:
Answer #1

Total estimated variable overhead=(3*30900)=$92700

Hence total estimated overhead=Total estimated variable overhead+Total estimated fixed overhead

=$92700+154500=$247200

Hence predetermined overhead rate=Total estimated overhead/Estimated total machine hours

(247200/30900)=$8 per machine hour

Hence applied overhead=($8 per machine hour*30 )

=$240.